OpenStudy (anonymous):

Integrate dx/sqrt(9 + x^2)

OpenStudy (anonymous):

looks like a hyperbolic sine

OpenStudy (nincompoop):

try trig substitution

OpenStudy (anonymous):

or this. \[\int\limits_{?}^{?}\frac{ dx }{ \sqrt{9+x ^{2}} } \]

OpenStudy (nincompoop):

yeah

OpenStudy (nincompoop):

@satellite73 actually gave out the clue I just realized that

OpenStudy (anonymous):

hmm. not 100 percent sure how to do this.

OpenStudy (anonymous):

\[ 9+9\tan^2(\theta)=9\sec^2(\theta) \]

OpenStudy (nincompoop):

there might be some other ways, but quite cumbersome

OpenStudy (anonymous):

Consider \(x^2=9\tan^2(\theta)\) and \(dx=3\sec^2(\theta)d\theta\)

OpenStudy (nincompoop):

I don't think he's familiar with trig substitution yet...

OpenStudy (anonymous):

why does x^2 equal that though? We had went over it a little my last day of class but then we had spring break so i didnt quite grasp it.

OpenStudy (nincompoop):

http://www.math.smith.edu/Local/cicintro/ch11.pdf

OpenStudy (anonymous):

Well, can you think of any other nice trig identity where \(1+f^2(x) = g^2(x)\)?

OpenStudy (anonymous):

You could also try \(9 + 9\cot^2(\theta) = 9\csc^2(\theta)\) as well if you want. Then \(x^2=9\cot^2(\theta)\).

OpenStudy (anonymous):

so... are we using trig functions just because we could set any value we wanted for x? im not quite sure what the reasoning is for this trig to come in which is why i am lost :(

zepdrix (zepdrix):

@jeremyggg The reason for the Trig Sub is that making use of Trig Identities allows us to get rid of addition/subtraction signs between terms. If I have something like this:\[\Large\bf\sf \sqrt{1+\tan^2 x}\]I can't take the square root easily, the addition is causing problems! :( But if we make use of our Trig Identity,\[\Large\bf\sf \sqrt{1+\tan^2x}\quad=\quad \sqrt{\sec^2x}\]It becomes a lot more manageable!\[\Large\bf\sf =\quad \sec x\]Does that help explain it a lil bit? Trig Identities make our integrands much easier to work with.

OpenStudy (anonymous):

it helps but i dont understand where we are turning \[\sqrt{9+x ^{2}}\] into a trig identity?

zepdrix (zepdrix):

All of the square identities involve 1 + (something)^2. See how we have a 9 and not a 1? Bad bad bad, we'll have to get it into the form 1+(something)^2. Factoring a 9 out of each term, \[\Large\bf\sf \sqrt{9+x^2}\quad=\quad \sqrt{9\left(1+\frac{1}{9}x^2\right)}\]We'll do some sneaky magic and bring the 1/9 into the square as (1/3)^2. (And also bring the 9 out of the root).\[\Large\bf\sf =\quad 3\sqrt{1+\left(\frac{x}{3}\right)^2}\] Now we have the form 1+(something)^2 under the root. The Square Identity Involving addition is \(\Large\bf\sf 1+\tan^2\theta=\sec^2\theta\). So we're going to make the substitution, \[\Large\bf\sf \frac{x}{3}\quad=\quad \tan \theta\]

zepdrix (zepdrix):

\[\Large\bf\sf 3\sqrt{1+\left(\frac{x}{3}\right)^2}\quad=\quad 3\sqrt{1+\left(\tan \theta\right)^2}\quad=\quad 3\sqrt{1+\tan^2\theta}\]We can make use of our Trig Identity from there, yes? :o

OpenStudy (anonymous):

0.o yes!

OpenStudy (anonymous):

hmm. so i got \[\int\limits_{?}^{?}\frac{ dx }{ \sqrt{9(1+\tan ^{2}(\theta)} }\] and then \[\frac{ 1 }{ 3 }\int\limits_{?}^{?}\frac{ dx }{ \sqrt{\sec ^{2}\theta} }\] is this the right track?

zepdrix (zepdrix):

Yes very good. It can simplify a bit further right? You can take the square root of a square. Then we need to switch out the differential for something involving \(\Large d\theta\).

zepdrix (zepdrix):

Before we can integrate*

zepdrix (zepdrix):

Trig Substitutions are a long process. 1. Getting everything in terms of theta is a bit of work. 2. Simplifying and then integrating isn't too bad (since you've made it easier to work with). 3. Undoing your substitution is a bit of a pain. You have to draw out a triangle and use your trig relations (Soh Cah Toa) to get your result back in terms of x. I can explain it if you get there at some point :) Seems like you're still working on steps 1 and 2 though for now, which is fine.

OpenStudy (anonymous):

the dx is \[\sec ^{2}\theta d\theta\] so im at \[\frac{ 1 }{ 3 }\int\limits_{?}^{?}\sec\theta\ d \theta \]

OpenStudy (anonymous):

do i mess with the back sub before i integrate?

OpenStudy (anonymous):

1/3 lncos(theta) +c